# references.bib

### v1

@@ Appendix A Factorized solutions: motivation
Here we motivate why if we start from two solutions $k^{R}$ and $\tilde{k}^{R}$ of the boundary Yang-Baxter equation for an R-matrix $r(u)$ of a spin 1/2 chain, then their tensor product is a solution in the enlarged space.
@@
We write the boundary YB equation [^16] For simplicity we omit the superscript $R$. (2.9) in the enlarged space, by making the identification $1=12$ and $2=34$. We obtain
@@
We can easily identify the two boundary YBE for each of the two spin 1/2 chains. This also explains why the free constants in $k^{R}(u)$ (we labelled them $\kappa$) and $\tilde{k}^{R}$ are independent. In other words, we proved that one solution of $K(u)$ can be obtained by taking the tensor product of two solutions of the boundary Yang-Baxter for the spin 1/2 chain.
@@ Appendix B Allowed transformations
It is known that given an $R$-matrix solution of the Yang-Baxter equation (2.1), certain transformations can be applied to it, ensuring it remains a solution. These transformations are described in details in [deLeeuw:2019vdb, classificationybandboost, paletta2023yang]. Here, we analyze what happens to the $K$-matrices if we apply these transformations. In what follow, we first write how each transformation acts on the $R$-matrix and then on the $K^{L/R}$. We omit the superscript $L/R$ since the transformations act on the same way on both the $K$-matrices.
@@ Appendix C Explicit calculation of the first two derivatives of $t(u)$
We start from the expression of the transfer matrix with all the inhomogeneities set to 00 and we give the analytical expressions of the higher order derivatives. We use the expression (2.11) for the double raw transfer matrix. By omitting the dependence on the spectral parameter, this is
@@
where $A_{i}(u)$ are the matrices entering in the transfer matrix definition. We remark that this is also the assumption used by Sklyanin in [Sklyanin:1987bi] and that, for the case of the XX spin chain with magnetic field $h=1$ that we treat in section 5.2.3, the assumption does not hold.
@@
For simplicity, in this appendix we refer to $q^{(2)}_{i,j}=H_{i,j}$ as the range $2$ conserved density. For the models treated in the main text, this is the superoperator $\mathcal{L}$ given in (5.8).
@@ C.2 Second derivative $t^{\prime\prime}(0)$
We did not find this explicit computation in the literature. In [cuerno1993free], the authors gave the expression of $t^{\prime\prime}(u)$ for the case with $\text{tr}_{0}K_{0}^{L}(0)H_{L,0}\propto\mathbb{I}$ and $\text{tr}_{0}K_{0}^{L}(0)=0$. We checked that, under these assumptions, our expressions reduce to theirs. In that case, the expression of the second derivative, takes the expression (C.6). However, in the absence of these assumptions, the charge cannot be written in a form similar to (C.6). The terms acting in the boundary will have interaction range $2$ and the ones in the bulk range $3$.
[…]

### v2

@@ Appendix A Factorized solutions: motivation
Here we motivate why if we start from two solutions $k^{R}$ and $\tilde{k}^{R}$ of the boundary Yang-Baxter equation (10) for an R-matrix $r(u)$ of a spin 1/2 chain, then their tensor product is a solution in the enlarged space.
@@
We write the boundary YB equation [^20] For simplicity we omit the superscript $R$ for the right $K$-matrix. (10) in the enlarged space, by making the identification $1=12$ and $2=34$. We obtain
@@
We can easily recognize the two boundary YBE for each of the two spin 1/2 chains. This also explains why the free constants in $k^{R}(u)$ (we labelled them $\kappa$) and $\tilde{k}^{R}$ (labelled as $\tilde{\kappa}$) are independent. In other words, we proved that one trivial solution of $K(u)$ in the enlarged space can be obtained by taking the tensor product of two solutions of the boundary Yang-Baxter for the spin 1/2 chain.
@@ Appendix B Allowed transformations
It is known that given an $R$-matrix solution of the Yang-Baxter equation (2), certain transformations can be applied to it, ensuring it remains a solution. These transformations are described in details in [50, 51, 52]. Here, we analyze what happens to the $K$-matrices if we apply these transformations. In what follow, we first write how each transformation acts on the $R$-matrix and then on the $K^{L/R}$. We omit the superscript $L/R$ since the transformations act on the same way on both the $K$-matrices.
@@ Appendix C Explicit calculation of the first two derivatives of $t(u)$
We start from the expression of the transfer matrix with all the inhomogeneities set to 00 and we give the analytical expressions of the higher order derivatives. We use the expression (12) for the double raw transfer matrix. By omitting the dependence on the spectral parameter, this is
@@
where $A_{i}(u)$ are the matrices entering in the transfer matrix definition. We remark that this is also the assumption used by Sklyanin in [45] and that, for the case of the XX spin chain with magnetic field $h=1$ that we treat in section 5.2.3, the assumption does not hold. We motivate it in section C.1.
@@
For simplicity, in this appendix we refer to $q^{(2)}_{i,j}=h_{i,j}$ as the range $2$ conserved density. For the models treated in the main text, this is the superoperator $\mathcal{L}$ given in (105).
@@ C.1 First derivative $t^{\prime}(0)$
For the XX spin chain with magnetic field $h=1$ that we treat in section 5.2.3, the first derivative of the transfer matrix generates the dynamics. For that model, an interesting phenomenon happens and the expression (C.4) does not hold. In fact, the term $2\text{tr}_{0}K_{0}^{L}(0)h_{L0}$ that comes from taking the derivative of $R_{0L}(u)$ ($\text{lim}_{u\to 0}\text{tr}_{0}K_{0}^{L}(u)R_{0L}^{\prime}(u)R_{0,L-1}(u)\dots$) is divergent. However, adding it to the terms coming from $\text{lim}_{u\to 0}\text{tr}_{0}{K_{0}^{L}}^{\prime}(u)R_{0L}(u)R_{0,L-1}(u)\dots$ cancels the divergent terms. Since (C.3) holds, we may be tempted to conclude that for the term added, the following identity holds
$\text{lim}_{u\to 0}\text{tr}_{0}{K_{0}^{L}}^{\prime}(u)R_{0L}(u)R_{0,L-1}(u)\dots=\text{lim}_{u\to 0}\text{tr}_{0}{K_{0}^{L}}^{\prime}(u).$ (C.8)
This is not correct for this model, in fact the r.h.s. is divergent, while the l.h.s. is finite. To obtain the correct result, one should first add the terms coming from the derivatives of $R_{0L}$ and $K_{0}^{L}$ and only afterward take the limit.
@@ C.2 Second derivative $t^{\prime\prime}(0)$
We did not find this explicit computation in the literature. In [53], the authors gave the expression of $t^{\prime\prime}(u)$ for the case with $\text{tr}_{0}K_{0}^{L}(0)h_{L,0}\propto\mathbb{I}$ and $\text{tr}_{0}K_{0}^{L}(0)=0$. We checked that, under these assumptions, our expressions reduce to theirs. In that case, the expression of the second derivative, takes the expression (C.6). However, in the absence of these assumptions, the charge cannot be written in a form similar to (C.6). The terms acting in the boundary will have interaction range $2$ and the ones in the bulk range $3$.